# House of Cards: Massive Weights in LLMs

### v1

@@ Abstract
Massive activations, which manifest in specific feature dimensions of hidden states, introduce a significant bias in large language models (LLMs), leading to an overemphasis on the corresponding token. In this paper, we identify that massive activations originate not from the hidden state but from the intermediate state of a feed-forward network module in an early layer. Expanding on the previous observation that massive activations occur only in specific feature dimensions, we dive deep into the weights that cause massive activations. Specifically, we define top-$k$ massive weights as the weights that contribute to the dimensions with the top-$k$ magnitudes in the intermediate state. When these massive weights are set to zero, the functionality of LLMs is entirely disrupted. However, when all weights except for massive weights are set to zero, it results in a relatively minor performance drop, even though a much larger number of weights are set to zero. This implies that during the pre-training process, learning is dominantly focused on massive weights. Building on this observation, we propose a simple plug-and-play method called MacDrop (massive weights curriculum dropout), to rely less on massive weights during parameter-efficient fine-tuning. This method applies dropout to the pre-trained massive weights, starting with a high dropout probability and gradually decreasing it as fine-tuning progresses. Through experiments, we demonstrate that MacDrop generally improves performance across zero-shot downstream tasks and generation tasks.
@@ 1 Introduction
In this paper, we first delve deeper into massive activations, providing two key observations. (1) The bos token always has massive activations in the same feature dimensions and makes attention sinks, regardless of its position in the sequence. This observation enables us to focus only on the feature dimensions, instead of both the sequence and feature dimensions, when addressing massive activations. Namely, a simplified and consistent analysis of massive activations can be achieved by using only the bos token. (2) Massive activations originate in the intermediate state $\hat{h}^{inter}_{l}$ within an early layer $l$, before appearing in the hidden state $h_{l}$, as illustrated in Figure 1(a). Namely, massive activations triggered in $\hat{h}^{inter}_{l}$ are subsequently and continuously propagated through skip connections. This observation implies that the feed-forward network in layer $l$ plays a crucial role in LLMs.
Finally, we propose a straightforward plug-and-play method during parameter-efficient fine-tuning, named massive weights curriculum dropout (MacDrop). This method applies dropout to the pre-trained massive weights, rather than additional trainable weights, starting with a high dropout rate that is progressively reduced throughout the fine-tuning phase. The intuition behind MacDrop is that a high initial dropout rate encourages the model to lessen dependence on the massive weights predominantly learned during the pre-training phase. Then, reducing the dropout rate facilitates a more stable convergence, ensuring the pre-trained model is leveraged with neglectable damage by the end of fine-tuning. In zero-shot downstream tasks and generation tasks, we demonstrate that MacDrop generally enhances model performance.
@@ 2 Massive Weights
In this section, we review the key observations on massive activations reported by Sun et al. (2024) and extend the analysis by exploring various states using the bos token, which was not covered. Based on this expanded analysis, we formally define top-$k$ massive weights in a specific layer and investigate their importance through two opposite types of attacks.
@@ 2.1 Prerequisite: Massive Activations
$h_{l}=\hat{h}_{l}+\text{FFN}(\text{LN}(\hat{h}_{l})),\text{where}\,\hat{h}_{l}=h_{l-1}+\text{ATTN}(\text{LN}(h_{l-1}))\vspace*{-5pt}$ (1)
Sun et al. (2024) primarily concentrated on the activations within hidden states, identifying that certain activations exhibit exceptionally large magnitudes, which they termed massive activations. Massive activations are observed at the starting position (i.e., input-agnostic) or at the delimiter tokens, depending on the model. Furthermore, these activations are confined to a small number of fixed feature dimensions, even within these tokens. These activations initially emerge after passing through several early layers and then decreases as they near the last layer.
Massive activations are strongly tied to the attention sinks phenomenon, as identified by Xiao et al. (2024), in which attention is abnormally concentrated on a small subset of tokens. In detail, a given query state tends to have positive cosine similarity with the key states of the tokens exhibiting massive activations, and negative cosine similarity with those of other tokens. Consequently, attention is heavily skewed toward the tokens associated with massive activations.
@@ 2.2 Further Analysis on Massive Acitvations
bos token always has massive activations, regardless of its poisition.
Figure 2: (Top) Magnitudes of the hidden state and (Bottom) attention scores after Softmax, according to the position of the bos token. The described hidden state is the output of layer 2 (i.e., $h_{2}$). The attention scores are calculated at layer 3 (i.e., after massive activations appear) and averaged across different heads. The bos token always has massive activations, regardless of its position.
We begin by examining whether any specific token consistently triggers massive activations. As noted by Sun et al. (2024), massive activations are observed when any token is placed in the starting position; however, no single token has been found to trigger massive activations regardless of its position. The existence of such a token would greatly facilitate the analysis and algorithm development for handling massive activations, making the process more deterministic by focusing on that token.
Based on the previous observation that nonsemantic tokens can trigger massive activations in certain LLMs, we focus our analysis on the bos token. Figure 2 describes the magnitudes of activations of the hidden state and normalized attention scores according to the position of the bos token, after massive activations appear. We use the implementation [^1] https://github.com/locuslab/massive-activations of massive activations for this example and visualization. Llama-3-8B does not have massive activations at delimiter tokens such as ‘.’ (first column). When the bos token is placed in the starting position (second column), the ‘Summer’ token loses its massive activations, while when the bos token is placed in the middle or ending position (third and fourth columns), it triggers massive activations in the same feature dimensions (i.e., 788, 1384, and 4062). What is intriguing is that, despite the difference in magnitude according to the position, the bos token similarly exhibits attention sinks. Therefore, we use only the bos token (placed at the starting position) for the continuation of analysis and algorithm development.
(a) Various states in early layers.
(b) Hidden states.
(c) Intermediate states.
Figure 3: (a) Magnitudes of various states and (b and c) the top three and median magnitudes of hidden states and intermediate states across layers. These results show that massive activations in the hidden state originate from those in the interemediate state in a FFN module in an early layer.
Furthermore, we decompose a FFN module into ${\bm{W}}_{down}(\sigma({\bm{W}}_{gate}(\cdot))\odot{\bm{W}}_{up}(\cdot))$, to analyze the intermediate states (i.e., the output of $\sigma({\bm{W}}_{gate}(\cdot))\odot{\bm{W}}_{up}(\cdot)$). Figure 3(c) describes the top three and median magnitudes of the intermediate state across layers. It is demonstrated that massive activations originate in the intermediate state of a FFN module in an early layer $l$. This result implies that ${\bm{W}}_{up}$ and ${\bm{W}}_{gate}$ in layer $l$ are closely tied to massive activations. Additional results for other LLMs are provided in the Appendix C.
2.3 Massive Weights
@@
To assess the significance of massive weights, we conduct two types of attacks: top-$k$ zeroing and top-$k$ retaining. Note that these attacks only affect the ${\bm{W}}_{up}$ and ${\bm{W}}_{gate}$ projection matrices in layer $l$, where massive weights are present. The first attack is to set the top-$k$ massive weights to zero (i.e., darker orange weights in Figure 1(a)). In essence, this attack is very similar to the one proposed in Sun et al. (2024), where massive activations in the hidden state are zeroed out in a single layer. The difference is that their attack targets the hidden state, while our attack targets the intermediate state. The second attack is to set all weights to zero except for top-$k$ massive weights (i.e., lighter orange weights in Figure 1(a)). That is, the number of rows being damaged in each attack is $k$ and the dimensions of intermediate state -- $k$, respectively.
Following Sun et al. (2024), we assess perplexity11footnotemark: 1 on three datasets: WikiText (Merity et al., 2017), C4 (Raffel et al., 2020), and PG-19 (Rae et al., 2020). Additionally, we evaluate zero-shot accuracy [^2] https://github.com/EleutherAI/lm-evaluation-harness on five tasks: Arc-Easy, Arc-Challenge (Clark et al., 2018), BoolQ (Clark et al., 2019), PIQA (Bisk et al., 2020), and WinoGrande (WG) (Sakaguchi et al., 2021). Table 1 presents the results of two attacks on Llama-3-8B, Llama-3-70B, and Llama-3.1-405B (8bit) [^3] We use 8xA100-80G GPUs for our work; therefore, we employ an 8-bit model., when $k$ is set to $5$. Llama-3-8B has its massive weights in layer 2 out of 32 layers, Llama-3-70B in layer 4 out of 80 layers, and Llama-3.1-405B in layer 6 out of 126 layers (Appendix B). The larger the difference from the original performance, the stronger the attack.
Top-5 zeroing is a much stronger attack than top-5 retaining, even though top-5 retaining sets several thousands times more weights to zero compared to top-5 zeroing does for the same projection matrices. This means that in the projections ${\bm{W}}_{up}$ and ${\bm{W}}_{gate}$ at layer $l$, having only massive weights is significantly better than having all weights except for massive weights. In detail, similar to the findings of Sun et al. (2024), the top-$k$ zeroing attack proves to be highly effective in disrupting the Llama-3 family, even for extremely large-scale models (e.g., 70B and 405B). On the other hand, the top-$k$ retaining attack does not cause complete damage. In conclusion, these findings reveal that massive weights are predominantly learned during pre-training, highlighting their essential contribution to the overall performance of LLMs.
$k$, which affects performance, depends on the model architecture.
@@
The Llama families are highly sensitive to massive weights. In the top-$k$ zeroing, a noticeable performance drop occurs even when $k$ is as small as 3, irrespective of the model’s scale. In top-$k$ retaining, when $k$ is set to 1 (i.e., with only one row active in ${\bm{W}}_{up}$ and ${\bm{W}}_{gate}$ in layer $l$), the performance nearly reaches the original level in smaller-scaled models ($\leq$ 13B). While, for larger-scaled models ($\geq$ 70B), the top-30 massive weights are required to maintain performance.
Similarly, Mistral is also disrupted, when $k$ is set to 5. Mixtral is a sparse Mixture of Experts (MoE) architecture that uses a top-2 routing mechanism, where two experts are activated among eight FFN modules in each layer. To attack the Mixtral model, we identify the active experts in the layer with massive weights using the bos token. Figure 5 describes the probability distribution of experts in the Mixtral model across all layers. Notably, it is observed that when massive activations occur, a single expert (i.e., expert 4) is assigned a significantly higher probability than the others. Therefore, we target only the ${\bm{W}}_{up}$ and ${\bm{W}}_{gate}$ of this expert, rather than all experts. Although Mixtral does not completely break down, there is a considerable decline in performance when the top-50 massive weights are zeroed out. These results indicate that, despite the immense resources required to build high-performance LLMs, they can collapse like a house of cards even under minimal attacks.
@@
$h_{l}=\hat{h}_{l}+\text{dropout}(\text{FFN}(\text{LN}(\hat{h}_{l}))),\text{where}\,\hat{h}_{l}=h_{l-1}+\text{dropout}(\text{ATTN}(\text{LN}(h_{l-1})))$ (2)
The Gemma-2 family is exceptionally resilient to the top-$k$ zeroing attack, maintaining almost no loss in performance even when $k$ is large. Additionally, even if ${\bm{W}}_{up}$ and ${\bm{W}}_{gate}$ are entirely eliminated (i.e., $k=0$ in top-$k$ retaining), there is no noticeable performance degradation. This family incorporates two additional LN layers after both the ATTN and FFN modules, formed by Eq. (3). These added normalization layers result in completely different hidden and intermediate states compared to other models, as described in Appendix C. Furthermore, attention sinks at the initial token are not observed in the Gemma-2 family (Appendix D).
$h_{l}=\hat{h}_{l}+\text{LN}(\text{FFN}(\text{LN}(\hat{h}_{l}))),\text{where}\,\hat{h}_{l}=h_{l-1}+\text{LN}(\text{ATTN}(\text{LN}(h_{l-1})))$ (3)
Nevertheless, because most existing open-source models, except for a few, are still heavily dependent on massive weights, it is critical to take them into account.
@@ 3 Massive Weights Curriculum Dropout
In this section, we propose a straightforward plug-and-play method, termed massive weights curriculum dropout (MacDrop), during parameter-efficient fine-tuning such as low rank adaptation (Hu et al., 2022). This method applies dropout to the pre-trained massive weights with a curriculum that gradually reduces the dropout probability. The reason for applying dropout to weights (Wan et al., 2013) instead of activations (Srivastava et al., 2014) is that the number of massive activations is only $k$, but that of massive weights is $k\times d$, where $d$ is the dimension of the hidden states. Note that our method is not applied to the trainable parameters of adapters. Therefore, MacDrop can be applied orthogonally to the process of training the adapter. Algorithm 1 describes MacDrop in a pseudo PyTorch (Paszke et al., 2019) style, and is implemented within the trainer code of transformers [^4] https://github.com/huggingface/transformers/tree/main/src/transformers. Initially, massive weights are identified using the bos token before fine-tuning (Lines 1-3). Subsequently, an adapter is trained while the pre-trained massive weights are dropped. Meanwhile, a curriculum strategy is applied to progressively enable the use of the original pre-trained weights without masking. Note that in Line 8, ‘model’ includes both the masked pre-trained network and a trainable adapter.
MacDrop is motivated by the observation that massive weights are predominantly learned during the pre-training phase, and that zeroing them can severely undermine LLMs. Therefore, at the early stages of fine-tuning, the objective is to reduce the reliance on massive weights, as their excessive dominance may lead the model to over-rely on specific patterns. Moreover, considering that the undamaged pre-trained model is used after fine-tuning is finished, we develop a strategy to adjust the dropout rate using a curriculum.
// Dropout is only executed in layer $l$,
// where the massive intermediate state $h^{inter}_{l}$ appears.
@@
1 _, sorted_indices = torch.sort(torch.abs($h^{inter}_{l}$), descending=True)
2 massive_indices = sorted_indices[:$k$]
// massive weights
3 massive_W_up = copy(W_up[massive_indices, :]); massive_W_gate = copy(W_gate[massive_indices, :])
4 for $t=1$ to $T$ do
       // decreasing dropout probability
5       $p$ = $p_{0}\times(1-\frac{t}{T})$
       // pre-trained massive weights dropout
6       mask = (torch.rand(massive_W_up.shape) $>$ $p$).int()
7       W_up[massive_indices, :] *= mask; W_gate[massive_indices, :] *= mask
8       tr_loss_step = training_step(model, inputs)
       // pre-trained massive weights rollback
9       W_up[massive_indices, :] = massive_W_up; W_gate[massive_indices, :] = massive_W_gate
10
Algorithm 1 Top-$k$ Massive Weights Curriculum Dropout in pseudo PyTorch style
4 Experiments
4.1 Zero-shot Downstream Task
@@
Model Method ARC-Easy ARC-Challenge BoolQ PIQA WinoGrande Avg.

### v2

@@ Abstract
Massive activations, which manifest in specific feature dimensions of hidden states, introduce a significant bias in large language models (LLMs), leading to an overemphasis on the corresponding token. In this paper, we identify that massive activations originate not from the hidden state but from the intermediate state of a feed-forward network module in an early layer. Expanding on the previous observation that massive activations occur only in specific feature dimensions, we dive deep into the weights that cause massive activations. Specifically, we define top-$k$ massive weights as the weights that contribute to the dimensions with the top-$k$ magnitudes in the intermediate state. When these massive weights are set to zero, the functionality of LLMs is entirely disrupted. However, when all weights except for massive weights are set to zero, it results in a relatively minor performance drop, even though a much larger number of weights are set to zero. This implies that during the pre-training process, learning is dominantly focused on massive weights. Building on this observation, we propose a simple plug-and-play method called MacDrop (massive weights curriculum dropout), to rely less on massive weights during parameter-efficient fine-tuning. This method applies dropout to the pre-trained massive weights, starting with a high dropout probability and gradually decreasing it as fine-tuning progresses. Through various experiments, including zero-shot downstream tasks, long-context tasks, and ablation studies, we demonstrate that MacDrop generally improves performance and strengthens robustness.
Machine Learning, ICML
@@ 1 Introduction
In this paper, we first delve deeper into massive activations, providing two key observations. (1) The bos token placed at the starting position always has massive activations in the same feature dimensions and makes attention sinks. This observation enables us to focus only on the feature dimensions, instead of both the sequence and feature dimensions, when addressing massive activations. Namely, a simplified and consistent analysis of massive activations can be achieved by using only the bos token. (2) Massive activations originate in the intermediate state $\hat{h}^{inter}_{l}$ within an early layer $l$, before appearing in the hidden state $h_{l}$, as illustrated in Figure 1(a). Namely, massive activations triggered in $\hat{h}^{inter}_{l}$ are subsequently and continuously propagated through skip connections. This observation implies that the feed-forward network in layer $l$ plays a crucial role in LLMs.
Finally, we propose a straightforward plug-and-play method during parameter-efficient fine-tuning, named massive weights curriculum dropout (MacDrop). This method applies dropout to the pre-trained massive weights, rather than additional trainable weights, starting with a high dropout rate that is progressively reduced throughout the fine-tuning phase. The intuition behind MacDrop is that a high initial dropout rate encourages the model to lessen dependence on the massive weights predominantly learned during the pre-training phase. Then, reducing the dropout rate facilitates a more stable convergence, ensuring the pre-trained model is leveraged with neglectable damage by the end of fine-tuning. Through various ablation studies, we examine the effects of dropout scope and dropout probability scheduling. Finally, we demonstrate that MacDrop generally enhances model performance and robustness in zero-shot downstream tasks and long context tasks.
@@ 2 Massive Weights
In this section, we review the key observations on massive activations reported by Sun et al. (2024a) and extend the analysis by exploring various states using the bos token, which was not covered thoroughly. Based on this expanded analysis, we formally define top-$k$ massive weights in a specific layer and investigate their importance through two opposite types of attacks.
@@ 2.1 Prerequisite: Massive Activations
$\begin{split}h_{l}&=\hat{h}_{l}+\text{FFN}(\text{LN}(\hat{h}_{l})),\\ \text{where}\,\hat{h}_{l}&=h_{l-1}+\text{ATTN}(\text{LN}(h_{l-1}))\end{split}$ (1)
Sun et al. (2024a) primarily concentrated on the activations within hidden states, identifying that certain activations exhibit exceptionally large magnitudes, which they termed massive activations. Massive activations are observed at the starting position (i.e., input-agnostic) or at the delimiter tokens, depending on the model. Furthermore, these activations are confined to a small number of fixed feature dimensions, even within these tokens. These activations initially emerge after passing through several early layers and then decreases as they near the last layer.
Massive activations are strongly tied to the attention sinks phenomenon, as identified by Xiao et al. (2024), in which attention is abnormally concentrated on a small subset of tokens. In detail, a given query state tends to have positive cosine similarity with the key states of the tokens exhibiting massive activations, and negative cosine similarity with those of other tokens. Consequently, attention is heavily skewed toward the tokens associated with massive activations. Detailed related work is explained in Appendix H.
Figure 2: (Top) Magnitudes of the hidden state and (Bottom) attention scores after Softmax of Mistral-7B, according to the position of the bos token. The described hidden state is the output of layer 16 (i.e., $h_{16}$). The attention scores are calculated at layer 17 (i.e., after massive activations appear) and averaged across different heads.
@@ 2.2 Further Analysis on Massive Acitvations
bos token placed at the starting position always has massive activations.
We begin by examining whether any specific condition consistently triggers massive activations. The existence of such a condition would greatly facilitate the analysis and algorithm development for handling massive activations. Following Sun et al. (2024a), massive activations are observed when any token is placed at the starting position; however, we find the cases where the token at the starting position does not trigger massive activations and attention sinks, such as Mistral-7B.
@@
Figure 2 describes the magnitudes of activations of the hidden state and normalized attention scores of Mistral-7B according to the position of the bos token, after massive activations appear. The reason for arbitrarily inserting the bos token is based on the previous observation that nonsemantic tokens can trigger massive activations in certain LLMs. We use the implementation [^1] https://github.com/locuslab/massive-activations of massive activations for this example and visualization. Mistral-7B has massive activations at the first delimiter token ‘.’, not at the starting position (first column). However, when the bos token is placed at the starting position, it triggers massive activations and the first delimiter token loses its massive activations (second column). When the bos token is inserted in the middle or ending position after the first delimiter token, massive activations are observed in both tokens (third and fourth columns). On the other hand, there are models that respond to the starting position but not to the bos token, such as Llama-2, detailed in Appendix C. Therefore, by considering both conditions, we use only the bos token placed at the starting position for the continuation of analysis and algorithm development.
(a) Various states in early layers.
(b) Hidden states.
(c) Intermediate states.
Figure 3: (a) Magnitudes of various states and (b and c) the top three and median magnitudes of hidden states and intermediate states across layers. These results show that massive activations in the hidden state originate from those in the intermediate state in a FFN module in an early layer.
Furthermore, we decompose a FFN module into ${\bm{W}}_{down}(\sigma({\bm{W}}_{gate}(\cdot))\odot{\bm{W}}_{up}(\cdot))$, to analyze the intermediate states (i.e., the output of $\sigma({\bm{W}}_{gate}(\cdot))\odot{\bm{W}}_{up}(\cdot)$). Figure 3(c) describes the top three and median magnitudes of the intermediate state across layers. It is demonstrated that massive activations originate in the intermediate state of a FFN module in an early layer $l$. This result implies that ${\bm{W}}_{up}$ and ${\bm{W}}_{gate}$ in layer $l$ are closely tied to massive activations. Additional results for other LLMs are provided in the Appendix D.
2.3 Massive Weights
To assess the significance of massive weights, we conduct two types of attacks: top-$k$ zeroing and top-$k$ retaining. Note that these attacks only affect the ${\bm{W}}_{up}$ and ${\bm{W}}_{gate}$ projection matrices in layer $l$, where massive weights are present. The first attack is to set the top-$k$ massive weights to zero (i.e., darker orange weights in Figure 1(a)). In essence, this attack is very similar to the one proposed in Sun et al. (2024a), where massive activations in the hidden state are zeroed out in a single layer. The difference is that their attack targets the hidden state, while our attack targets the intermediate state. The second attack is to set all weights to zero except for top-$k$ massive weights (i.e., lighter orange weights in Figure 1(a)). That is, the number of rows being damaged in each attack is $k$ and the dimensions of intermediate state -- $k$, respectively.
@@ 2.3 Massive Weights
Following Sun et al. (2024a), we assess perplexity11footnotemark: 1 on three datasets: WikiText (Merity et al., 2017), C4 (Raffel et al., 2020), and PG-19 (Rae et al., 2020). Additionally, we evaluate zero-shot accuracy [^2] https://github.com/EleutherAI/lm-evaluation-harness on five tasks: Arc-Easy (ARC-E), Arc-Challenge (ARC-C) (Clark et al., 2018), BoolQ (Clark et al., 2019), PIQA (Bisk et al., 2020), and WinoGrande (WG) (Sakaguchi et al., 2021). Table 1 presents the results of two attacks on Llama-3-8B, Llama-3-70B, and Llama-3.1-405B (8bit) [^3] We use 8xA100-80G GPUs for our work; therefore, we employ an 8-bit model., when $k$ is set to $5$. Llama-3-8B has its massive weights in layer 2 out of 32 layers, Llama-3-70B in layer 4 out of 80 layers, and Llama-3.1-405B in layer 6 out of 126 layers. Detailed indices of massive weights are provided in Appendix B, including various LLMs. The larger the difference from the original performance, the stronger the attack.
Top-5 zeroing is a much stronger attack than top-5 retaining, even though top-5 retaining sets several thousands times more weights to zero compared to top-5 zeroing does for the same projection matrices. This means that in the projections ${\bm{W}}_{up}$ and ${\bm{W}}_{gate}$ at layer $l$, having only massive weights is significantly better than having all weights except for massive weights. In detail, similar to the findings of Sun et al. (2024a), the top-$k$ zeroing attack proves to be highly effective in disrupting the Llama-3 family, even for extremely large-scale models (e.g., 70B and 405B). On the other hand, the top-$k$ retaining attack does not cause complete damage. In conclusion, these findings reveal that massive weights are predominantly learned during pre-training, highlighting their essential contribution to the performance of LLMs.
The value of $k$, which affects performance, varies depending on the model architecture.
@@
The Llama families are highly sensitive to massive weights. In the top-$k$ zeroing, a noticeable performance drop occurs even when $k$ is as small as 3, irrespective of the model’s scale. In top-$k$ retaining, when $k$ is set to 1 (i.e., with only one row active in ${\bm{W}}_{up}$ and ${\bm{W}}_{gate}$ in layer $l$), the performance nearly reaches the original level in smaller-scaled models ($\leq$ 13B). While, for larger-scaled models ($\geq$ 70B), the top-30 massive weights are required to maintain performance. Moreover, since different versions of Llama families pre-trained on varying datasets (e.g., with different cutoff dates) exhibit similar tendencies, it can be inferred that the pre-training dataset itself does not significantly influence the emergence of massive phenomena.
Similarly, Mistral is also significantly disrupted, when $k$ is set to 5. Mixtral is a sparse Mixture of Experts (MoE) architecture that uses a top-2 routing mechanism, where two experts are activated among eight FFN modules in each layer. To attack the Mixtral model, we identify the active experts in the layer with massive weights using the bos token. Figure 5 describes the probability distribution of experts in the Mixtral model across all layers. Notably, it is observed that when massive activations occur, a single expert (i.e., expert 4) is assigned a significantly higher probability than the others. Therefore, we target only the ${\bm{W}}_{up}$ and ${\bm{W}}_{gate}$ of this expert, rather than all experts. Although Mixtral does not completely break down, there is a considerable decline in performance when the top-50 massive weights are zeroed out. These results indicate that, despite the immense resources required to build high-performance LLMs, they can collapse like a house of cards even under minimal attacks.
@@
$\begin{split}h_{l}&=\hat{h}_{l}+\text{dropout}(\text{FFN}(\text{LN}(\hat{h}_{l}))),\\ \text{where}\,\hat{h}_{l}&=h_{l-1}+\text{dropout}(\text{ATTN}(\text{LN}(h_{l-1})))\end{split}$ (2)
The Gemma-2 family is exceptionally resilient to the top-$k$ zeroing attack, maintaining almost no loss in performance even when $k$ is large. Additionally, even if ${\bm{W}}_{up}$ and ${\bm{W}}_{gate}$ are entirely eliminated (i.e., $k=0$ in top-$k$ retaining), there is no noticeable performance degradation. This family incorporates two additional LN layers after both the ATTN and FFN modules, formed by Eq. (3). These added normalization layers result in completely different hidden and intermediate states compared to other LLMs, as described in Appendix D. Furthermore, attention sinks are not observed in the Gemma-2 family, as shown in Appendix E.
$\begin{split}h_{l}&=\hat{h}_{l}+\text{LN}(\text{FFN}(\text{LN}(\hat{h}_{l}))),\\ \text{where}\,\hat{h}_{l}&=h_{l-1}+\text{LN}(\text{ATTN}(\text{LN}(h_{l-1})))\end{split}$ (3)
These observations imply that LLMs not sensitive to massive phenomena may not perform effectively with algorithms based on massive phenomena. Further discussion is provided in Appendix F.
@@ 3 Massive Weights Curriculum Dropout
In this section, we propose a straightforward plug-and-play method, termed massive weights curriculum dropout (MacDrop), during parameter-efficient fine-tuning such as low rank adaptation (Hu et al., 2022). This method applies dropout to the pre-trained massive weights with a curriculum that gradually reduces the dropout probability. The reason for applying dropout to weights (Wan et al., 2013) instead of activations (Srivastava et al., 2014) is that the number of massive activations is only $k$, but that of massive weights is $k\times d$, where $d$ is the dimension of the hidden states. It is important to note that our method does not modify the trainable parameters of adapters; instead, it is applied to the pre-trained frozen weights. Therefore, MacDrop can be applied orthogonally to the process of training the adapter.
MacDrop is motivated by the observation that massive weights are predominantly learned during pre-training, and that zeroing them can severely undermine LLMs. Therefore, at the early stages of fine-tuning, the objective is to reduce the reliance on massive weights, as their excessive dominance may lead the model to over-rely on specific patterns. Moreover, considering that the undamaged pre-trained model is used after fine-tuning is finished, we develop a strategy to adjust the dropout rate using a curriculum.
Algorithm 1 describes MacDrop in a pseudo PyTorch (Paszke et al., 2019) style, and is implemented within the trainer code of transformers [^4] https://github.com/huggingface/transformers/tree/main/src/ transformers. Initially, massive weights are identified using the bos token before fine-tuning (Lines 1-4). Subsequently, an adapter is trained while the pre-trained massive weights are dropped. Meanwhile, a curriculum strategy is applied to progressively enable the use of the original pre-trained weights without masking. Note that in Line 10, ‘model’ includes both the masked pre-trained network and a trainable adapter. When we implement Lines 1-4 in practice, we precomputed the massive indices and loaded them. Lines 6-9 and 11-12 are the additional computation for MacDrop, which requires neglectable overhead (e.g., approximately 0.35 second per step for Llama-3-8B using LoRA on 8xA100 GPUs).
Algorithm 1 Top-$k$ massive weights curriculum dropout (MacDrop) in pseudo PyTorch style
// Dropout is only executed in layer $l$, where $h^{inter}_{l}$ appears.
@@
_, sorted_indices = torch.sort(torch.abs($h^{inter}_{l}$), descending=True)
massive_indices = sorted_indices[:$k$]
// pre-trained massive weights
massive_W_up = copy(W_up[massive_indices, :])
massive_W_gate = copy(W_gate[massive_indices, :])
for $t=1$ to $T$ do
2       // decreasing dropout probability $p$ = $p_{0}\times(1-\frac{t}{T})$
// pre-trained massive weights dropout mask = (torch.rand(massive_W_up.shape) $>$ $p$).int()
W_up[massive_indices, :] *= mask
W_gate[massive_indices, :] *= mask
tr_loss_step = training_step(model, inputs)
// pre-trained massive weights rollback W_up[massive_indices, :] = massive_W_up
W_gate[massive_indices, :] = massive_W_gate
Table 2: Long context tasks performance.
Single-document QA Multi-document QA Summarization Few-shot learning Synthetic Code
Method
NrtvQA
Qasper
MF-en
HotpotQA
2WikiMQA
Musique
GovReport
QMSum
MultiNews
TREC
TriviaQA
SAMSum
PCount
PRe
Lcc
RB-P
Avg.
Model: Meta-Llama-3-8B
LoRA 26.03 30.38 53.38 26.30 23.05 11.96 29.00 22.81 26.43 72.50 81.14 44.27 2.63 32.00 72.90 69.70 39.03
+ MacDrop 25.31 34.05 46.84 38.06 28.99 17.92 29.62 22.86 26.64 72.00 89.34 45.08 3.00 27.50 73.17 68.25 40.54
DoRA 26.31 31.57 52.10 27.04 23.57 12.01 29.20 23.35 26.34 73.50 81.35 43.22 2.61 30.00 73.46 69.44 39.07
+ MacDrop 26.11 30.99 53.37 29.35 25.66 12.14 29.07 23.10 26.32 73.50 86.63 44.64 2.05 25.50 73.90 69.32 39.48
Model: Mistral-7B-v0.1
LoRA 22.52 34.64 35.65 32.11 19.80 12.73 27.22 21.98 26.73 69.00 87.50 41.70 1.00 21.00 71.57 65.47 36.91
+ MacDrop 23.49 38.51 36.11 37.78 27.60 14.91 26.40 22.53 26.92 69.50 89.92 37.07 1.55 20.00 70.37 65.87 38.03
DoRA 22.79 34.52 35.55 30.87 17.84 12.26 27.45 22.15 26.52 70.00 88.05 41.56 1.00 20.50 71.88 65.38 36.77
+ MacDrop 23.10 35.10 35.00 29.53 23.77 10.50 27.14 22.63 27.50 69.00 89.56 38.96 1.00 21.42 71.17 65.53 36.93
@@
Model Method ARC-E ARC-C BoolQ PIQA WG Avg.